OpenStudy (anonymous):

find the integral of 1/x^2(sqrt(4-x^2) using trig substitution

OpenStudy (anonymous):

I did try I ended up with a really complicated integral since I also substituted the 4sin^2theta for the x

OpenStudy (anonymous):

2cos tdt=dx

OpenStudy (anonymous):

4sin^2t=x^2

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Wait did you ask me to try to do this function? @OOOPS ???

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

x = 2sec(t) dx = 2sec(t)tan(t) Integral ( [1/sqrt(4sec^2(t) - 4)] [2sec(t)tan(t)] )

OpenStudy (anonymous):

Wait am I doing it right?

OpenStudy (anonymous):

Integral ( [2sec(t)tan(t)] / [2tan(t)] )dt Integral ( sec(t) ) dt

OpenStudy (anonymous):

= ln |sec(t) + tan(t)| + C

OpenStudy (anonymous):

That is all that I know how to do.

OpenStudy (kainui):

I think you should be using:\[x=2 \sin \theta\]

OpenStudy (anonymous):

Umm I'm getting confused by all the replies a bit

OpenStudy (anonymous):

Sorry.

OpenStudy (anonymous):

Yes I made x that however x^2 confuses me since I end up with 1/4sin^2theta after the cancellations and I cannot past that

OpenStudy (anonymous):

*get past I mean

OpenStudy (kainui):

The idea behind trig substitution is that you can turn something like: \[\sqrt{1-\cos^2 \theta}\] into \[\sqrt{\sin^2 \theta}=\sin \theta\] because \[\sin^2 \theta + \cos^2 \theta = 1\] and for your problem you have to change it by multiplying the whole thing by 4. \[4\sin^2 \theta + 4 \cos^2 \theta = 4\]\[4-4\sin^2 \theta = 4 \cos^2 \theta\]

OpenStudy (anonymous):

yes I get that under the square root and cancel it with the 2costheta dtheta generated in the numerator however the 4sin^2theta leftover for the x^2 in the denominator

OpenStudy (kainui):

Consider the following: \[\frac{ 1 }{ \sin \theta }=\csc \theta\]

OpenStudy (anonymous):

I don't understand how will this help me

OpenStudy (kainui):

Try plugging it in, maybe I'll remind you of one other thing... But not quite what you need... but almost... \[\frac{ d }{ dx } \tan \theta = \sec^2 \theta\]

OpenStudy (kainui):

I guess I should have written: \[\frac{ d }{ d \theta}\] but you know.

OpenStudy (anonymous):

I'm sorry I'm still a little confused

OpenStudy (anonymous):

I would get integral csc^2theta which is cottheta this is still pretty complicated

OpenStudy (anonymous):

cot tan theta

OpenStudy (kainui):

\[-\cot \theta = \int\limits_{}^{}\csc^2 \theta d \theta\] That's pretty much all you need, not too complicated.

OpenStudy (anonymous):

That isn't what the answer key says though it says -\[-\sqrt{4-x ^{2)}}/4x\]

myininaya (myininaya):

You have to finish in terms of x of course.

OpenStudy (kainui):

I'll walk through the basic: \[x=2\sin \theta\] will give you:\[\int\limits_{}^{}\frac{ 2\cos \theta d \theta }{ 4\sin^2 \theta *2 \cos \theta }=\int\limits_{}^{}\csc^2\theta d \theta\]\[=-\frac{ 1 }{ 4 } \cot \theta\] Now you need to plug back in for x!

myininaya (myininaya):

Use the substitution you made.

OpenStudy (anonymous):

So x/2 is sintheta and dx/2 is cos theta

OpenStudy (anonymous):

and cottantheta is cos/sin

OpenStudy (kainui):

\[\frac{ -1 }{ 4 }\cot(\sin^{-1}(\frac{ x }{ 2 }))\] That looks sloppy though, and besides it's simpler if we just think back to this: \[\frac{ x }{ 2}=\sin \theta\] and now we remember SOH CAH TOA and draw this out:  Now when you do this:: \[\frac{ -1 }{4 }\cot(\theta)\] just use that triangle, remember cot is just adjacent over opposite.

OpenStudy (kainui):

And oh, hey, it should match what your answer is. =)

OpenStudy (anonymous):

Thank you that matches the answer. :)

OpenStudy (kainui):

Glad I could help, I love trig substitution along with pretty much all of calculus. =)